DOUBT OF SEQUENCE AND SERIES....
an=2n-5/2 ;FIND : t8,t12

Grade:12

3 Answers

Muskaan
7 years ago
An and tn means the same thing T8 = 2^8-5/2=507/2,T12= 2^12 - 5/2= 16379/2Therefore , answer is 507/2, 16379/2
VIKRAM ENIMIREDDY
7 years ago
an=2^n -5/2
a8=2^8-5/2
   =(2^9-5)/2
   =(512-5)/2

1. a8 =507/2
a12=2^12-5/2
    =(2^13-5)/2
   =(8192-5)/2     

2. a12=8187/2
now i think you can satisfied with this format.
Shubham Pitroda
7 years ago
No...VIKRAM ENIMIREDDY.....tn and an is different form....tn is nth term and by an we can find the first term and second term.......Solution:first put 1 and 2 in an simultaneously where a1 is first term and a2 is second term... By this we can find the , difference also.....by this we can find t8 and t12 also.......This is the perfect solution........